Find the slope-intercept form for line line that satisfy the conditions given. slope 5, passing through (1,2).
Answer by srinivas.g(540)

slope intercept form : y= mx+c
where m =slope
c= y-intercept
slope (m)=5
hence y= 5x+c-------------------------eq(1)
this line is passing through (1,2)
hence, put x,y values in the above equation
2= 5*1+c
2=5+c
move 5 to the left
2-5=c
-3= c
c=-3
put c= -3 in eq(1)
Hence required equation y= 5x-3